OpenStudy (anonymous):

Given that int_{1}^{3} f(x) dx = -20 and int_{1}^{3} [f(x) + kx^2] dx = 6 find the value of k

OpenStudy (hoblos):

\[\int\limits(n + m)dx = \int\limits n dx + \int\limits mdx \]

OpenStudy (hoblos):

so here \[\int\limits_{1}^{3} [f(x) + kx^2] dx = \int\limits_{1}^{3} f(x)dx + \int\limits_{1}^{3} kx^2 dx\]

OpenStudy (hoblos):

do you know how to calculate \[\int\limits_{1}^{3} kx^2 dx \] ?

OpenStudy (anonymous):

maybe yes

OpenStudy (hoblos):

so try to solve it and let me see what you get

OpenStudy (anonymous):

ok, i will try.

OpenStudy (anonymous):

 is it correct so far?